OpenStudy (anonymous):

Help?? Subtract the polynomials, showing all the steps ( 9x - 4) – (3x + 2) = 2. ( 3x^2 + 8) – (2X^2 + 1 ) =

OpenStudy (sleepyjess):

(9x-4)-(3x+2) the - in front of the second polynomial means you will multiply (3x+2) by -1 can you do that?

OpenStudy (sleepyjess):

@loopgirl I need you to at least try before I can move forward

OpenStudy (anonymous):

Sorry, but i don't understand.

OpenStudy (sleepyjess):

what is 3*-1?

OpenStudy (anonymous):

I don't understand where you got -1

OpenStudy (sleepyjess):

from the - in front of 3x+2

OpenStudy (anonymous):

Ok

OpenStudy (sleepyjess):

so all that - means is that you are going to change the signs in the parentheses so - becomes + and + becomes - 3*-1 becomes -3 so what do you think +2 will be?

OpenStudy (anonymous):

-2

OpenStudy (sleepyjess):

correct so now we can remove the parentheses from the entire equation now we have 9x-4-3x-2 now combine like terms and you have the answer

OpenStudy (anonymous):

so -6?

OpenStudy (sleepyjess):

what is 9x-3x?

OpenStudy (anonymous):

6x

OpenStudy (sleepyjess):

correct and -4-2 is -6 which you already got so your answer will be 6x-6

OpenStudy (anonymous):

Ok, I understand now. thank you

OpenStudy (sleepyjess):

try the next one on your own and let me know what you come up with

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I got 17 - 5 but it doesn't seem right.

OpenStudy (sleepyjess):

what did you get after multiplying \((2x^2+1)*-1\)?

OpenStudy (anonymous):

I got -5

OpenStudy (sleepyjess):

what is \(2x^2*-1\)?

OpenStudy (anonymous):

4

OpenStudy (sleepyjess):

how are you getting 4? there is no value for x

OpenStudy (anonymous):

Exponent 2 * -1 = -2 -2 * -2 = 4

OpenStudy (sleepyjess):

no all you need to do is change the sign to begin with it is positive so now it will be negative

OpenStudy (anonymous):

what would the equation look like? because i would think that it would look like -2x^2 but i still don't think that is right.

OpenStudy (sleepyjess):

yes that is correct for the first part

OpenStudy (sleepyjess):

now what is 1*-1?

OpenStudy (anonymous):

-1

OpenStudy (sleepyjess):

correct so now your full equation is \(3x^2+8-2x^2-1\) now \(3x^2-2x^2=~\sf and~8-1\)

OpenStudy (anonymous):

would combine like terms?

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

so is this what it would look like? 1x^2 -7

OpenStudy (sleepyjess):

it would be x^2 but your last term is wrong should be 8-1

OpenStudy (anonymous):

OK thank you

OpenStudy (sleepyjess):

so your final answer will be \(x^2+7\)

OpenStudy (anonymous):

Ok